OpenStudy (anonymous):

how do you solve 4^x=20?

OpenStudy (jdoe0001):

well, I'd use the aforementioned logarith cancellation rule :)

OpenStudy (anonymous):

\[a^x=b\] \[x=\log _ab\]

OpenStudy (anonymous):

thanks you alot guys i apreciate it (:

OpenStudy (jdoe0001):

\(\bf \text{now using => }log_aa^x = x\\ 4^x = 20\\ log_44^x = log_420 \implies x = log_420\)

OpenStudy (jdoe0001):

then you'd use the "change of base rule" to get the value using log base 10

OpenStudy (anonymous):

that means i have to divide right?

OpenStudy (jdoe0001):

log change of base rule => \(\bf log_ab = \cfrac{log_{10}b}{log_{10}a}\)

OpenStudy (anonymous):

is it 2.16?

OpenStudy (jdoe0001):

that's what I got, yes

OpenStudy (campbell_st):

why not just use base e or base 10 logs\ \[x = \frac{\ln(20)}{\ln(4)}\] don't worry about change of base

OpenStudy (cwrw238):

or you could just take logs to base e of both sides ln 4^x = ln 20 and by the law of logs: x ln4 = ln 20 x = lm20 / ln 4